OpenStudy (anonymous):

omg, how do you evaluate an infinite series? lim To infinity, Sigma sign n=1 1 ----- n2 + n

OpenStudy (anonymous):

i think you have to compute partial sums yes?

OpenStudy (anonymous):

hmm, not sure

OpenStudy (anonymous):

I think I have to figure out what type of series it is first or some pellet

OpenStudy (anonymous):

i just did this in my summer class today let me get you a reference sheet my instructor gave to everyone it should help alot one moment

OpenStudy (anonymous):

wow holy pellet I love the internet!

OpenStudy (anonymous):

this site is unbelievable

OpenStudy (anonymous):

partial fractions i think gets it. let me see

OpenStudy (anonymous):

good work going to take a look now

OpenStudy (anonymous):

this converges by integral test, and in fact the sum is 1. but we need partial fractions i think to get the partial sums. i have to compute for a minute or several

OpenStudy (anonymous):

OH I get it now, partial fractions right

OpenStudy (anonymous):

for the part on e where it says "converges is |r| > 1" its suppose to say converges if |p| > 1

OpenStudy (anonymous):

how do you figure out if it is an "alternating series"

OpenStudy (anonymous):

riddle me this

OpenStudy (anonymous):

negative then to the power?

OpenStudy (anonymous):

ok yes if you use partial fractions you get it

OpenStudy (anonymous):

\[\frac{1}{n(n+1)}=\frac{1}{n}-\frac{1}{n+1}\] gives a telescoping sum

OpenStudy (anonymous):

keep going, and also, see my question above about 'alternating series'

OpenStudy (anonymous):

in fact you can instantly see the series converges to 1

OpenStudy (anonymous):

you take the limit of each term?

OpenStudy (anonymous):

I can't instantly see anything like that

OpenStudy (anonymous):

\[1-\frac{1}{2}+\frac{1}{2}-\frac{1}{3}+\frac{1}{3}-\frac{1}{4}+...=1\]

OpenStudy (anonymous):

it telescopes

OpenStudy (anonymous):

oh so it's sort of like limit pellet, you just need to recognize those patterns

OpenStudy (anonymous):

a, c d, e on the study guide are the only helpful things on it and i just started today on it so i hope satellite can clear the rest of it up XD

OpenStudy (anonymous):

yeah helps if you know the tricks

OpenStudy (anonymous):

even though it's 'converging to 1', doesn't it go to 0 since it starts at 1 and you are subtracting terms?

OpenStudy (anonymous):

bu you can see it yes? and if you want to be rigorous i guess using the telescoping nature you can actually write out what the partial sums look like.

OpenStudy (anonymous):

what are the "tricks"

OpenStudy (anonymous):

so some in here - Q4d, e, 5, 6

OpenStudy (anonymous):

the meaning of the infinite sum is the limit of the partial sums. the individual terms of course go to zero, otherwise you can forget it. but the question was not "does this converge" but rather "what is the limit?"

OpenStudy (anonymous):

thanks broseph

OpenStudy (anonymous):

you can actually write out the partial sums and see what they are. i mean write out \[\sum_{k=1}^n\frac{1}{k^2+k}\]

OpenStudy (anonymous):

ok hang on

OpenStudy (anonymous):

i believe if you do this, using the partial fraction decomposition, you will see that you get \[\sum_{k=1}^n\frac{1}{k^2+k}=\frac{n}{n+1}\] and then if you take \[\lim_{n\rightarrow \infty}\frac{n}{n+1}\] you get 1 by inspection

OpenStudy (anonymous):

forget that noise it is even easier

OpenStudy (anonymous):

ok so it's 1/k^2 + k = A/K + B/K+1 what do I do from here??

OpenStudy (anonymous):

\sum_{k=1}^n\frac{1}{k^2+k}=\frac{n}{n+1} where did you get that from?

OpenStudy (anonymous):

put \[S_n=\sum_{k=1}^n \frac{1}{k^2+k}=\sum_{k=1}^n\frac{1}{k}-\frac{1}{k+1}\]

OpenStudy (anonymous):

here we will see it. but actually easier.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so what do we have? \[S_1=1-\frac{1}{2}\] \[S_2=1-\frac{1}{3}\] \[S_3=1-\frac{1}{4}\] and in general by the telescope \[S_n=1-\frac{1}{n+1}\] which happens to be \[S_n=\frac{n}{n+1}\] but we don't really need that form

OpenStudy (anonymous):

the definition of the infinite sum is the limit as n goes to infinity of the partial sums

OpenStudy (anonymous):

and that limit is clear as day 1

OpenStudy (anonymous):

the whole gimmick was to use partial fractions and see that you ended up with the telescoping sum, making the partial sums easy to compute, and therefor the limit easy as well

OpenStudy (anonymous):

ok what are the tricks? which are the common ones to look for there are like 4

OpenStudy (anonymous):

n/n+1 is one of them for sure

OpenStudy (anonymous):

not sure what exactly you mean. sometimes the sums are impossible. sometimes you get \[\frac{\pi^2}{6}\] and some genius figured it out 200 years ago. but this partial sum one is a good one.

OpenStudy (anonymous):

i mean "partial fractions" one

OpenStudy (anonymous):

ok first step see if limit is equal to 0 that could mean it converges. lets see this baby

OpenStudy (anonymous):

you mean the limit of the terms, not the limit of the partial sums. the limit of the terms must be zero but you need lots more than that.

OpenStudy (anonymous):

right what else do you need?

OpenStudy (anonymous):

oh is that the one I'm supposed to be trying?

OpenStudy (anonymous):

i cant believe no one gave you a medal Satellite!